Find f'(x) of f(x)= sin[(sin⁷x⁷ + 1)⁷]


1
Expert's answer
2021-03-01T15:35:45-0500

Let us find "f'(x)" of "f(x)= \sin[(\sin^7x^7 + 1)^7]"


"f'(x)= \cos[(\sin^7x^7 + 1)^7]\cdot 7(\sin^7x^7 + 1)^6\cdot7\sin^6x^7\cdot \cos x^7\cdot 7x^6=243x^6 \cos[(\sin^7x^7 + 1)^7]\cdot (\sin^7x^7 + 1)^6\cdot\sin^6x^7\cdot \cos x^7"